Subject: APO 3.5-m Users Committee minutes 1/23/06
From: Bruce Gillespie
Submitted: Mon, 23 Jan 2006 16:14:03 -0700
Message number: 944
(previous: 943,
next: 945
up: Index)
APO 3.5-m Users Committee Phonecon, 1/23/06
Attending: Suzanne Hawley, Bruce Gillespie, Michael Strauss, Jon
Holtzman, John Bally, John Wilson, Bruce Balick, Al Harper, Russet
McMillan
Absent: Don York, Karl Glazebrook
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from Institutional representatives:
Princeton (Michael Strauss)--Michael said his people are pretty
happy, no complaints. He added that the PU users feel that there is
no problem with the current data security protocols. Data archiving
would be useful, but it shouldn't be a high priority for us.
Colorado (John Bally)--John reported that the CU users have no
concerns about data security, and an archive would be good but it's
not that important.
NMSU (John Holtzman)--Nothing significant to report.
Johns Hopkins--No report given.
Chicago (Al Harper)--nothing significant to report.
UWashington (Bruce Balick and Suzanne Hawley)--BruceB said he felt
the UW users were pretty contented. Suzanne added that a small
service-observing program has been successfully initiated for/with
George Wallerstein.
Univ. of Virginia (John Wilson)--John said he nothing to report.
**********************************
Discussion of telescope/instruments report:
BruceG summarized highlights from the report (see below). There was
no further discussion.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
11/23/05 through 1/19/06
Bruce Gillespie
0) Overview
Highlights include a good run of operability for the telescope and
instruments, a major design review for the Triplespec NIR
spectrograph project, and refinements to the 2006 plans for CIF
projects. NIC-FPS was put back into service, and the focus and stray
light problems in DIS were fixed. The unusually good weather for
this time of year continued, and was largely well used for observing.
1) Telescope
The 3.5-m telescope had very little in the way of problems during the
past two months. Part of one night was lost when a drive amplifier
blew a transistor, which was repaired quickly the next day. There
may be a software workaround to prevent this in the future, which is
being investigated.
2) The performance of the instruments was nominal, and noteworthy is:
o NIC-FPS: The vacuum-leak problem in the NIC-FPS dewar seems
to have been solved by the CU developers, and the instrument has been
used for science several times in the period. Also, the NIC-FPS
Fabry-Perot filter seems to have been repaired by the vendor, and we
are planning to test it cold in the instrument sometime in the near
future.
o DIS: The red camera focus was adjusted to good effect, and
the scattered light in the blue camera was determined to be some kind
of contamination inside the dewar, which was baked out while vacuum
pumping.
3) Engineering and CIF projects
A detailed plan and budget for the 2006 CIF projects was developed.
More progress was made on assembling the new drive boxes, which
should be finished next month. Design work continued on the new top
end, and materials have been received and machining work on some of
the parts has begun. Maintenance and enhancement of TUI and other
software systems continued. The JHU proposal to upgrade DIS was
turned into a work plan for 2006.
4) Miscellaneous
The 4th Observing Specialist position recruitment was completed;
Gabrelle Saurage, who has telescope operations experience both at HET
and Keck, has accepted our offer. She will begin training at APO
next month, and start work full time in early June.
APOLLO has been running for several partial nights each month, and is
semi-regularly getting lots of lunar-return photons.
**********************************
Triplespec review:
Suzanne reminded everyone that UVa is officially joining ARC in a
year, and their users are in the meantime getting familiarized with
APO through DD time and helping with CorMASS runs. Last month, there
was a final design review at APO for Triplespec, the NIR spectrograph
UVa is building as their capital buy-in to ARC. The formal review
team consisted of Alan Uomoto (chair), Steve Smee (JHU), Tom O'Brien
(OSU), Craig Loomis (PU), and Al Harper (UC). The review
demonstrated that the UVa instrument components were largely coming
along well, and John Wilson said that since the review, the
components being built by Cornell and Caltech were building
momentum. All in all, the review went well, and an excellent review
committee report was submitted to UVa. We expect UVa to respond to
the reviewers' comments by mid-February. As part of the UVa/ARC
agreement, we are purchasing the Rockwell detector using CIF monies;
the order is being placed now and we should have the chip in ~6 mos.
or less, with the finished instrument on the telescope in about a
year. There is an issue with the NA2 rotator and Triplespec (weight
and balance) that is being addressed by a joint APO/UVa engineering
collaboration.
**********************************
Futures Committee:
Suzanne reported that the ARC Futures Committee held a meeting at JHU
last month, and after 1-1/2 years of work, the group is converging on
a final report. Nearly all of the background material has been
assembled, and a suite of models of how to run the observatory (both
telescopes) after 2008 is being prepared. A draft of the report is
expected to be sent to the Futures Committee members in early
February, and they will distribute the report to their respective
constituencies for feedback. After assessing the community feedback,
the final report will be submitted to the ARC Board of Governors in
the Mar/Apr time frame. Based on the recommendations of the report,
ARC will probably have just enough time to implement various
programmatic options to put us on a stable footing for continued
observatory operations in the post-SDSS-II era.
ACTION: The Users Committee members will need to help disseminate
the draft report to their users.
**********************************
Observing software and data access:
Suzanne said that Remark is now unavailable for remote observing--
we've pulled the plug, so to speak. Russet said that nobody has used
it remotely since last summer. BruceG mentioned that Russell Owen
wants to ask all Mac TUI users to upgrade to OS version 10.4. There
was general discussion that brought out that this may be inconvenient
and/or difficult for some users.
ACTION: BruceG will ask Russell what the pros and cons are of
keeping TUI compatible with older versions of operating systems and
libraries.
ACTION: Users Committee members should poll their users to ask if
there are compelling reasons to keep TUI operable without using the
latest version of operating systems.
On the data access issue, Suzanne said that we have no budget set
aside to operate and maintain a data archive for the 3.5-m telescope,
but we now have enough disk space to allow data to remain intact at
APO for from 6 to 9 months after taken. We still recommend that
people download their data immediately after observing, because the
on-site storage is not backed up. The concern over data security was
briefly discussed, and it was decided that since nearly all the users
believe their data are secure enough now, if specific concerns about
proposal and data security are brought to the observatory's
attention, they can be handled on a case-by-case ad hoc basis.
**********************************
[open from previous months]:
ACTION: Once we collect the the names of the designated
institutional systems administrators, BruceG will have them listed on
the APO web pages for reference. STATUS: will be completed in January.
ACTION: BruceG will write a paragraph about the importance of having
back-up programs, to be added to the e-mail automated message we send
observers prior to their runs. STATUS: will be completed in January.
[new actions from last month]:
ACTION: Users Committee members should poll their users to ask if
the the present data distribution/security systems at APO are
adequate, and if not, why. STATUS: Discussed at January telecon.
ACTION: Users Committee members should poll their users to ask if we
should permanently archive all 3.5-m data, and possibly make it
available to the public, and if so, why. STATUS: Discussed at
January telecon.
ACTION: Suzanne and BruceG will discuss this [the DIS red focus
problem] in the upcoming weekly operations telecon. STATUS:
Completed. DIS cameras were refocussed in December by Russet
McMillan and Jon Holtzman--much improvement. We are planning a DIS-
focus monitoring program for the future, one that routinely
compensates for seasonal changes.
**********************************
Extra topic-DIS upgrade report:
Jon Holtzman said that the JHU contract has been issued for new
gratings and the blue-prism upgrade. The procurement of the new red
deep-depletion detector has been delayed by the vendor; we're
probably looking at summertime installations of all these upgrades.
**********************************
Next Users Committee phonecon will be on Monday, 27 February, at 8:30
AM Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 944 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO